OpenStudy (anonymous):

The new business building uptown is building a wall made entirely of glass pyramids. Each pyramid will have a square base with length of 8 inches and height of 10 inches. If the wall requires 250 of these pyramids, what volume of glass will be used? Round to the nearest cubic inch.

OpenStudy (anonymous):

volume=a^2*h/3 so v=(64)*(64)*(10)/3

OpenStudy (anonymous):

so solve that @Aliypop remeber / means divide

OpenStudy (anonymous):

13,653.333333

OpenStudy (anonymous):

yes so round to the nearest inch

OpenStudy (anonymous):

13,653

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help. @Aliypop